OpenStudy (appleduardo):

whats the integral of ln(x^(2) + 1)??

OpenStudy (appleduardo):

so far I have:\[\ln(x ^{2}+1)x-\int\limits_{}^{}x*\frac{ 2x }{ x ^{2}+1 } dx\]

sam (.sam.):

Long division for the int. by parts?

OpenStudy (appleduardo):

uhmm could that work?

OpenStudy (anonymous):

my guess is parts

OpenStudy (anonymous):

oh nvm you have that already

sam (.sam.):

\[\ln(x ^{2}+1)x-\int\limits\limits_{}^{}2-\frac{2}{x^2+1} dx \]

OpenStudy (anonymous):

\[\frac{x^2}{x^2+1}=\frac{x^2+1}{x^2+1}-\frac{1}{x^2+1}=1-\frac{1}{x^2+1}\]

OpenStudy (anonymous):

first integral is obvious, second part is the derivative of arctangent

OpenStudy (anonymous):

i ignored the 2, @.Sam. has it

OpenStudy (anonymous):

you could divide too, but i find that gimmick easier than long division

OpenStudy (appleduardo):

thank you! I'll try and post it when im done :P